OpenStudy (anonymous):

(2x10 to fifth power) x (4 x 10 to the 4th power) The answer that I received is 80,000,000.00 Is this correct?

OpenStudy (anonymous):

not correct. decimal point should be removed and the 0's after kept :P \[x^a * x^b = x^{a+b}\] \[2 * 10^5 \times 4*10^4 = (2*4)\times 10^{5 + 4} = 8 * 10^9\]

OpenStudy (chmvijay):

noo it must be 8,000,000,000.00

OpenStudy (anonymous):

no it musn't.. @chmvijay . only if it were 2.00e5 * 4.00e4

OpenStudy (chmvijay):

nope it can 8*10^9 =8*1000000000=8,000,000,000 or 2*10^5 =200000 and 4*10^4 =40000 =200000 * 40000 =8000,000,000 ur getting confusion with addition where it cantbe posbile

OpenStudy (anonymous):

9 decimal points, right? I often get confused where to place the decimal.

OpenStudy (chmvijay):

after 9 you can place or not to place it never mint its like $=50 or $=50.00 it wont make much different